OpenStudy (anonymous):

Choose one of the factors of x3 + 343. x - 7 x2 - 14x + 49 x2 + 7x + 49 None of the above

OpenStudy (anonymous):

@SarahBrown @kaylarocks001 @mathman806 @ejune420 @CamReLind

OpenStudy (jdoe0001):

keep in mind that \(\bf 343 = 7^3\\ \quad \\ a^3+b^3 = (a+b)(a^2-ab+b^2)\)